OpenStudy (anonymous):

The Earth is approximately 1.5*10^8km from the sun, and light travels at 3*10^5 km/s. How many seconds does light take to travel from the sun to the Earth?

OpenStudy (anonymous):

\[(1.5\times10^8\text{ km})\times\frac{1\text{ sec}}{3\times10^5\text{ km}}=0.5\times10^{3}\text{ sec}\]